Determine the solutions to each of the following independent cases: (a.) Collings College has annual fixed operating costs of $12,500,000 and variable operating costs of $1,000 per student. Tuition is $8,000 per student for the coming academic year, with a projected enrollment of 1,500 students. Expected revenues from endowments and federal and state grants total $250,000. Determine the amount the college must obtain from other sources. $Answer

(b.) The Collings College Student Association is planning a fall concert. Expected costs (renting a hall, hiring a band, etc.) are $30,000. Assuming 3,000 people attend the concert, determine the break-even price per ticket. How much will the association lose if this price is charged and only 2,700 tickets are sold? (Do not use a negative sign with your answer.) $Answer

(c.) City Hospital has a contract with the city to provide indigent health care on an outpatient basis for $25 per visit. The patient will pay $5 of this amount, with the city paying the balance ($20). Determine the amount the city will pay if the hospital has 10,000 patient visits. $Answer

(d.) A civic organization is engaged in a fund-raising program. On Civic Sunday, it will sell newspapers at $1.25 each. The organization will pay $0.75 for each newspaper. Costs of the necessary permits, signs, and so forth are $500. Determine the amount the organization will raise if it sells 5,000 newspapers. $Answer

(e.) Christmas for the Needy is a civic organization that provides Christmas presents to disadvantaged children. The annual costs of this activity are $5,000, plus $10 per present. Determine the number of presents the organization can provide with $20,000. Answer presents
Answer #1

a)

Annual fixed operating costs $12,500,000
Annual Variable operating costs (1500 * 1000) $1,500,000
Total costs 14,000,000
Less: Tuition revenue (8000*1500) -12,000,000
Less: Federal adn state grants -250,000
Amount obtained from other sources 11,750,000

b) Break Even Price per ticket = Total fixed costs / expeced tickets to be sold = $30000 / 3000 = $10 per ticket

Loss (in case 2700 tickets sold) = $30000 - $(2700* 10 ) = $3000

c) Out of $25 conract health care charges, the city pays $20.

Thus, The amount paid by city, if 10000 patient visits = $20 * 10000 = $200000

d) If 5000 newspapers wer sold, the net amount raised by the organisation = total revenue - total variable costs - total fixed costs = (5000 * $1.25) - (5000 * $0.75) - $500 = $6250 - $3750 - $500 = $2000

e) Number of presents = (Total amount provided - Fixed cost ) / cost incurred per present = ($20000 - $5000) / $10 = 1500 presents